There are two types of tickets available for a concert. The reserved-seat tickets cost $18.00 each, and the cheap-seat tickets cost $12.50 each. How many tickets were sold if 3 times as many cheap-seat tickets were sold as reserved-seat tickets and the total proceeds were $688,200?

What's the formula that's the thing I don't understand with story problems.

Answer by checkley71(8403)

12.5*3X+18*X=688,200
37.5+18X=688,200
55.5X=688,200
X=688,200/55.5
X=12,400 RESEVED SEAT TICKETS WERE SOLD.
3*12,400=37,200 OF THE CHEAP TICKETS WERE SOLD.
PROOF:
12.5*37,200+18*12,400=688,200
465,000+223,200=688,200
688,200=688,200